OpenStudy (anonymous):

My teacher gave an example of converting a function into a logarithmic function and it is.. f(x)=106.16e^(-0.73x) I understood the example up until In y - In 106.16 = In e^(-0.73x) But then the next step was In y = In e^(-0.73x) + 4.66 I just have no clue where the 4.66 came from? He didn't explain the steps, just laid it out.

zepdrix (zepdrix):

\[\large f(x)=106.16e^{(-0.73x)}\]Let f(x) = y\[\large y=106.16e^{(-0.73x)}\]Take the natural log of both sides.\[\large \ln y = \ln(106.16e^{(-0.73x)})\]

zepdrix (zepdrix):

\[\large \ln y = \ln106.16+\ln(e^{(-0.73x)})\]

zepdrix (zepdrix):

Looks like teacher simplified the first term on the right. \[\large \ln 106.16=4.66494739...\]

zepdrix (zepdrix):

\[\large \ln y = \ln(e^{(-0.73x)})+4.66\]

zepdrix (zepdrix):

The logarithmic and exponential are INVERSE operations of one another. So since we are taking the natural log of an exponential e, the essentially "Cancel out". \[\large \ln(e^{(-0.73x)})=(-0.73x)\]

zepdrix (zepdrix):

\[\large \ln y = -0.73x + 4.66\]